OpenStudy (jaimestibbard):

simplify using only positive exponents: 3.) 2( 2/2-x) [ -2/(2-x)^2]

zepdrix (zepdrix):

Hey Jaime! Welcome to OpenStudy! It's a little difficult to read your expression. Is this what you meant?\[\large\rm 2\left(\frac{2}{2-x}\right)\left[\frac{-2}{(2-x)^2}\right]\]

zepdrix (zepdrix):

If so, let's multiply the numerators together and the denominators with one another as well. (Treat the lonely 2 in front as a numerator value.\[\large\rm \frac{2(2)(-2)}{(2-x)(2-x)^2}\]Pull the negative out front for the multiplication,\[\large\rm \frac{-(2)(2)(2)}{(2-x)(2-x)^2}\]From there it should be pretty straight forward... just count your factors. Example:\[\large\rm (5)(5)(5)(5)=5^4\]

zepdrix (zepdrix):

Another example:\[\large\rm (x+y)(x+y)(x+y)=(x+y)^3\]